# A Coordinated Routing Approach for Enhancing Bus Timeliness and Travel Efficiency in Mixed-Traffic Environment

### v2

@@ IV-B Route Optimization for Rerouted CAVs
To mitigate traffic congestion on edge $(v_{b,k},v_{b,k+1})$ while improving the traffic efficiency of the transportation network, CAVs identified for rerouting are each assigned a new optimal route when $\hat{\tau}_{b,k}(t_{\text{sys}})\geq{\tau_{b,k}^{0}}$ is detected. Specifically, for each CAV $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, the system planner re-optimizes its route from intersection $v_{b,k}$ to the destination $d_{i}$, using the real-time traffic conditions observed at time $t_{\text{sys}}^{\prime}$. Based on the latest traffic data, including estimated travel times for each edge on GPLs and those on DLs, excluding the edge $(v_{b,k},v_{b,k+1})$, the optimization problem addressed for optimal route planning is cast as
@@
Constraints (5)-(8), (14), and (15) provide estimates of traffic flow and travel times for each edge on joint DLs, while constraints (9)-(12) offer travel time estimates for edges on GPLs. Moreover, (16) and (17) specify the triggering condition for rerouting and identify the set of CAVs selected for route replanning. After all estimated travel times across the network are updated at time $t_{\text{sys}}^{\prime}$, a prediction-aware Dijkstra algorithm [24] is executed to determine the optimal rerouting paths for all identified CAVs, denoted as $P_{i}^{*}(t_{\text{sys}}^{\prime})$, where $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$. In this modified Dijkstra framework, the edge weights correspond to the latest updated estimated travel times instead of static free-flow travel times or distances.
Remark 3
Solving the above route optimization problem requires real-time traffic data collected from existing sensor infrastructures, such as loop detectors and cameras. These data are aggregated within monitoring windows to compute vehicles’ actual travel times on the current edges and to estimate their travel times on future edges. Incorporating travel time estimations enables long-horizon congestion prediction and route adjustment in advance, thus preventing vehicles from being trapped in severe congestion.
@@ V Simulation Study
This section presents simulation studies conducted on the Simulation of Urban MObility (SUMO) platform to evaluate the effectiveness and overall performance of the proposed approach in a mixed traffic network. We start by introducing the simulation setup and parameter configurations, and then present simulation results and analysis. The code implementation is available online. [^1] See code at: https://github.com/Tanlu-L/Coordinated-Routing-Approach-for-Enhancing-Bus-Timeliness-and-Travel-Efficiency
@@
TABLE I: Scheduled bus arrival times
@@
V-A Simulation Setup
As shown in Fig. 4, we construct an urban road network in SUMO, which consists of $18$ road intersections and $3$ bus stops, with the road topology matching that shown in Fig. 1. This network replicates Van Ness Avenue in San Francisco from Market Street to Lombard Street, which includes one lane of DL. The roadway geometry and signal settings were obtained from OpenStreetMap. We consider that buses follow a predefined route along the joint DLs, traveling from the origin (node $7$ in Fig. 1) to the destination (node $15$). All other edges in the network are designated as GPLs. We assume sensors are installed at road intersections, enabling real-time observation of the number of vehicles entering each edge within a predefined monitoring time window. The monitoring time interval is set to $\Delta{T}\!=\!30$ seconds for joint DLs, and $\Delta\tilde{T}\!=\!60$ seconds for GPLs.
In the road network, we modeled the workday peak-hour bus service with $10$ buses operating along the same route from the origin stop to the destination stop, departing at intervals of $6$ minutes. The scheduled on-time arrivals at bus stops were determined based on free-flow travel times, with an additional $60$ seconds of dwell time at each station and a $30$-second disturbance margin to account for potential delays caused by traffic signals between stops. The scheduled arrival times for a representative bus are provided in Table I. CAVs and HVs enter the network at rates of $8$ and $20$ vehicles per minute, respectively, corresponding to a $30\%$ CAV penetration rate. The OD pair for each CAV is set as $(7,15)$, while the OD pair for each HV is randomly selected from the set $\{(1,6),(7,15),(16,21)\}$.
We compare the proposed coordinated routing strategy against two baseline methods: (i) static route planning (SRP) without rerouting and (ii) dynamic route planning (DRP) with rerouting. In the SRP strategy, routes are assigned based on free-flow travel times and remain fixed throughout the trip. In the DRP approach, each CAV is assigned the shortest path derived from real-time traffic data at the entry. Then, rerouting is applied if a lower-travel-time path is detected within a 60-second monitoring window. In the proposed method, CAVs follow the route recommendation obtained by the predictive control approach described in Section IV. Different from our approach, congestion anticipation on future road segments is not incorporated in the SRP and DRP. For both SRP and DRP, a modified Dijkstra algorithm using a weighted graph is applied to determine the shortest path. The weighted graph is initially constructed based on free-flow travel time on each edge, and the weights are updated dynamically according to the real-time traffic conditions.
Regarding the car-following-model in all three scenarios, CAVs use the intelligent driver model with a short headway ($0.6$ s) to capture automation-enabled efficiency. The maximum speed is set to $72\,\text{km/h}$, consistent with corridor speed limits. HVs and buses use the Krauss model: HVs with moderate headways (about $1.3$ s), variability to reflect human drivers, and a maximum speed of $72\,\text{km/h}$. Buses operate with larger headways ($1.6$ s) and a lower maximum speed of $65\,\text{km/h}$. This configuration reflects the capacity benefits of CAVs while realistically modeling human drivers and bus operations.
@@ V-B Results and Analysis
TABLE II: Bus on-time performance and delays at each station.
Station 1 Station 2 Station 3
On-time percentage [%]
   SRP using joint DL 10 30 30
   DRP using joint DL 50 60 60
   Proposed method 90 90 90
We evaluate the accumulated bus delay, defined as the total additional travel time experienced by all buses relative to their scheduled arrival times upon arrival at the destination, over a 3,600-second simulation horizon. The comparative results are presented in Fig. 5. As shown, buses experience significant delays under a static shortest-route assignment. The DRP approach reduces delays to some extent, but still, buses experience noticeable delays. In contrast, the proposed method significantly reduces bus delays nearly back to zero by integrating real-time traffic conditions and potential congestion estimation with efficient CAV rerouting. The results indicate that diverting CAVs away from congested edges, particularly those with bus stops during active dwell times, mitigates localized congestion along bus routes and improves schedule adherence. Table II reports the on-time percentage of $10$ buses at each station. The on-time percentage is defined as the fraction of buses arriving within $60$ seconds of their scheduled arrival times as given in Table I. It shows that the on-time percentages of the SRP, DRP, and our proposed method are approximately $23\%$, $57\%$, and $90\%$, respectively, demonstrating the superior performance of our method.
@@
Figure 8: The average bus delay at each stop.
Fig. 6 and Fig. 7 show the total travel time of CAVs and HVs over the simulation horizon. The proposed method achieves the lowest overall travel time for both CAVs and HVs. This improvement arises from proactively directing CAVs from congestion-prone edges, especially those affected by dwelling buses, thereby enhancing flow continuity throughout the network. Note that the DRP approach results in longer travel times for CAVs compared to both the SRP and the proposed method. Although for HVs, DRP initially reduces travel times, after $1800$ seconds, the travel times increase and even exceed those under SRP. These observations are common in transportation routing problems, as CAVs simultaneously identify and switch to shorter routes, which in turn leads to new and often more severe congestion on those routes.
Figure 9: The accumulated delay of each bus at experienced stops.
Fig. 8 shows the average trip delay for $10$ buses at each station, compared between the SRP method using the joint DL and the proposed approach. Bus stations indexed as $1$, 22, and $3$ correspond to nodes $9$, $11$, and $13$ in Fig. 1. As we can observe, the proposed method reduces bus delays significantly at all stations, achieving an average reduction of about $90.7\%$ per station compared with the SRP method. Fig. 9 further depicts the accumulated delay experienced by each bus across all stations along the route, showing that our rerouting strategy consistently reduces delays for individual buses at stop arrivals, thereby improving schedule adherence and overall transit reliability. Note that bus $6$ experiences relatively higher delays than others, mainly due to the great number of vehicles in the system. This result indicates that buses entering the network earlier or later experience less congestion. Such a pattern also aligns with the observed distribution of CAVs and HVs volumes in the system.
Figure 10: The average trip delay under different CAV penetration rates.
@@
Figure 11: 90th-percentile travel time comparison with and without joint DL.
To further evaluate the efficiency of the proposed method, we introduce a scenario termed SRP without using joint DL. In this scenario, we replicate the setting of a real dedicated bus lane by allowing only buses to use the lane, thereby ensuring that buses experience no delay, while CAVs follow their initially assigned shortest routes. The results show that allowing CAVs to share the joint DL reduces travel times for both CAVs and HVs, but at the expense of significantly increasing bus travel times. In contrast, the proposed method maintains bus travel times close to those in the dedicated bus lane scenario, while further reducing CAV and HV travel times compared to SRP with joint DL.
@@ VI Conclusions
In this paper, we presented a coordinated routing strategy to improve bus schedule adherence and enhance travel efficiency for CAVs in a mixed-traffic urban environment. A central element of our approach was a real-time traffic monitoring system that estimates traffic flow conditions on each road segment using sensor data. Leveraging these estimates, we developed a coordination scheme that identifies a minimal subset of CAVs for rerouting and assigns them optimized paths reflecting not only current but also anticipated potential congestion. By dynamically updating CAV routes in response to time-varying travel environments along DLs, our method effectively mitigates disruptions to transit operations. Simulation experiments conducted in SUMO based on realistic urban road network data demonstrate that the proposed strategy improves both transit reliability and CAV performance under diverse traffic scenarios. Future work will focus on extending this framework to accommodate multi-modal transportation systems and assessing its scalability in large-scale urban networks.

### v3

@@ IV-B Route Optimization for Rerouted CAVs
To mitigate traffic congestion on edge $(v_{b,k},v_{b,k+1})$ while improving the traffic efficiency of the transportation network, CAVs identified for rerouting are each assigned a new optimal route when $\hat{\tau}_{b,k}(t_{\text{sys}})\!\geq\!(1\!+\!\lambda){\tau_{b,k}^{0}}$ is detected. Specifically, for each CAV $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$, the system planner re-optimizes its route from intersection $v_{b,k}$ to $d_{i}$, using the real-time traffic conditions observed at time $t_{\text{sys}}^{\prime}$. Based on the latest traffic data, including the estimated travel times for each edge on GPLs and those on DLs, excluding the edge $(v_{b,k},v_{b,k+1})$, the optimization problem addressed for optimal route planning is cast as
@@
Constraints (4)-(7), (13), and (14) provide estimates of traffic flow and travel times for each edge on joint DLs, while constraints (8)-(11) offer travel time estimates for edges on GPLs. Moreover, (15) and (16) identify the set of CAVs selected for route replanning and specify the triggering condition for rerouting. After all estimated travel times across the network are updated at time $t_{\text{sys}}^{\prime}$, a prediction-aware Dijkstra algorithm [15] is executed to determine the optimal rerouting paths for all identified CAVs, denoted as $P_{i}^{*}(t_{\text{sys}}^{\prime})$, where $i\!\in\!\mathcal{R}_{b,k-1}(t_{\text{sys}}^{\prime})$. In this modified Dijkstra framework, the edge weights correspond to the latest updated estimated travel times instead of static free-flow travel times.
@@ V Simulation Study
This section presents simulation studies conducted on the SUMO platform to evaluate the effectiveness of the proposed approach. The code implementation is available online. [^1] See code at: https://github.com/Tanlu-L/Coordinated-Routing-Approach-for-Enhancing-Bus-Timeliness-and-Travel-Efficiency.git
V-A Simulation Setup
@@
As shown in Fig. 4, we construct an urban road network in SUMO with the road topology matching that shown in Fig. 1. This network replicates Van Ness Avenue in San Francisco from Market Street to Lombard Street. The roadway geometry and signal settings were obtained from OpenStreetMap. This corridor is representative of typical transit-priority urban arterials with DL operations. We consider that buses follow a predefined route along the joint DLs, traveling from the origin (node $7$ in Fig. 1) to the destination (node $15$). Sensors are assumed at road intersections, with monitoring windows set to $\Delta T\!=\!30$ s for joint DLs and $\Delta\tilde{T}\!=\!60$ s for GPLs.
TABLE I: Scheduled bus arrival times.
@@
We model a peak-hour service with $10$ buses departing every $6$ min from the same origin. The timetable in Table I is aligned with the real schedule and generated from free-flow travel times with $60$ s dwell per stop and a $30$ s signal margin. CAVs and HVs enter at rates of $8$ and $20$ veh/min, respectively ($30\%$ CAV penetration). The OD pair for each CAV is set as $(7,15)$, while the OD pair for each HV is randomly selected from the set $\{(1,6),(7,15),(16,21)\}$.
We compare the proposed coordinated routing strategy against two baseline methods: (i) static route planning (SRP) without rerouting and (ii) dynamic route planning (DRP) with rerouting. In the SRP strategy, routes are assigned based on free-flow travel times and remain fixed throughout the trip. In DRP, each CAV is initially assigned the shortest path using real-time traffic data and is rerouted if a lower-travel-time path is detected within a $60$ s monitoring window. In the proposed method, CAVs follow the predictive routing strategy described in Section IV. For both SRP and DRP, shortest paths are computed using a weighted Dijkstra algorithm, where edge weights are initialized by free-flow travel times and updated dynamically based on real-time traffic conditions.
@@ V-B Results and Analysis
TABLE II: Comparison of bus on-time percentage.
Methods Station 1 Station 2 Station 3
  SRP using joint DL $10\%$ $\% 3030\%$
  DRP using joint DL $50\%$ $\% 6060\%$
  Proposed method $\% 9090\%$ $90\%$
We evaluate the accumulated bus delay, defined as the total additional travel time experienced by all buses relative to their scheduled arrival times upon arrival at the destination, over a 3,600-second simulation horizon. The comparative results are presented in Fig. 5. As shown, buses experience significant delays under SRP, while DRP provides only partial improvement. In contrast, the proposed method nearly eliminates bus delays by integrating real-time traffic conditions and potential congestion estimation with efficient CAV rerouting. Table II reports the on-time percentage of the $10$ buses at each station, defined as arrivals within $60$ s of the scheduled times in Table I. The SRP, DRP, and proposed method achieve on average $23\%$, $57\%$, and $90\%$ on-time performance, respectively, demonstrating the advantage of the proposed approach.
Fig. 6 and Fig. 7 show the total travel time of CAVs and HVs over the simulation horizon. The proposed method achieves the lowest overall travel time for both vehicle classes by proactively diverting CAVs from congestion-prone edges, particularly those affected by bus dwell events, thereby improving network flow continuity. Note that the DRP approach yields longer CAV travel times than both SRP and the proposed method, a common effect in dynamic routing where vehicles simultaneously shift to shorter paths and induce secondary congestion.
@@
Figure 8: The trip delay in different CAV penetration.
@@
Figure 9: 90th-percentile travel time comparison.
To further evaluate the efficiency of the proposed method, we introduce a scenario termed SRP without joint DL access. In this setting, we emulate a dedicated bus lane by allowing only buses to use the DL, while CAVs follow their initially assigned shortest routes. Fig. 9 shows the 90th-percentile travel time for each vehicle type under SRP without joint DL, SRP with joint DL, and the proposed method. The comparison reveals that allowing CAVs to share the joint DL reduces travel times for both CAVs and HVs, but at the expense of increased bus travel time. In contrast, the proposed method maintains bus travel time close to that of the dedicated-lane scenario while further reducing CAV and HV travel times compared with SRP using joint DL.
@@ VI Conclusions
In this paper, we proposed a coordinated routing strategy to improve bus schedule adherence while enhancing CAV travel efficiency in mixed-traffic environments. The developed approach proactively identifies a minimal set of CAVs for rerouting based on real-time traffic monitoring conditions and anticipated potential congestion. Simulation experiments conducted in SUMO based on realistic urban road network data demonstrate that the proposed strategy improves both transit reliability and CAV performance across diverse traffic scenarios. A potential direction for future research is to focus on extending this framework to accommodate multi-modal transportation systems and assessing its scalability in large-scale urban road networks.